# Multi-language Unit Test Generation using LLMs

### v1

@@ 31                         $\mathcal{M}_{a}\leftarrow\mathcal{M}_{a}\cup{cs}$;
    Preprocessing for Python.  Unlike Java, Python’s more flexible structure, where a single file or module can contain a mix of functions, classes, and standalone code, lends itself better to module-scoped test generation. Therefore, for Python test generation, aster targets a module as a whole for test generation, including all classes, methods, and functions declared in the module. This module-level approach together with distinct language features of Python allows us to omit some preprocessing steps that are necessary for Java but irrelevant for Python. Because our approach provides the entire module to the LLM, instructing the LLMs about calls to private methods is unnecessary. Member visibility is Python is specified via naming conventions (names beginning with single underscore for protected members and double underscores for private members) and there is no strict encapsulation as these members can be accessed via name mangling. Python’s properties feature (using @property decorators) also lets private members to be accessed directly. Thus, identification of accessors is also unnecessary for Python test generation. In terms of relevant constructors, the constructor definitions in the focal module are already available to the LLM. Additionally, we add constructors for all imported modules to the LLM prompt. Also, we found that, for Python, adding a few examples of tests helps LLMs produce more predictable output, thus, making the postprocessing steps easier. For Java, this was not necessary, but investigation of RAG-based approaches for incorporating in-context learning for test generation is an interesting future research direction.
3.2. Postprocessing
3.2.1. Output sanitization
Generated tests undergo sanitization to remove extraneous content (such as natural language text) and ensure syntactic correctness of the generated test cases.
3.2.2. Error remediation
The sanitized code is compiled to identify any compilation errors. If errors occur, the error message as well as the compiler feedback is analyzed to identify the cause of the error. Section 3 of the LLM template updated with the erroneous line as well as the previous context to reprompt the LLM in a targeted manner to address the specified problems. Then, the compilable code is executed and runtime issues (assertion failure or runtime error) are fed to LLM for fixing. We use a similar prompt to provide the context related to the focal method and the error details.
3.2.3. Increasing coverage of generated tests
@@
Error-free test cases are executed to measure code coverage, which is further analyzed to identify uncovered code segments and then used to update the LLM prompt template (fig. 3) to guide the LLM to generate test cases that exercise the uncovered code segments. The process is repeated iteratively until the desired code coverage is achieved.
Postprocessing for Python. For Python, we use Pylint (pylint dev, [n. d.]), for identifying compilation and parsing error.
@@
Specifically, aster focuses on the error category of Pylint and uses this information to guide LLMs in the task of fixing compilation and parsing issues. Then, tests are executed using Pytest and the output is used to provide feedback to the LLMs in case of failures. Finally, Coverage.py (Batchelder, [n. d.]) is leveraged to perform the coverage-augmentation step.
Table 1. Models used in the evaluation.
@@
4. Experiment Setup
4.1. Research Questions
Our evaluation focuses on the following research questions:
RQ1::
@@ 4.1. Research Questions
RQ3::
How effective is aster in mutation scores achieved?
RQ4::
What is the effect of different types context information and postprocessing steps on coverage achieved?
RQ5::
Table 2. Java and Python datasets used in the evaluation.
@@ 4.2. Baseline test-generation tools
We evaluated aster against two state-of-the-art unit test generators. For Java, we used EvoSuite (evosuite, 2024), specifically Release 1.2.0. For Python, we selected CodaMosa and used the latest version from its repository (codamosa, 2024). CodaMosa is built on Pynguin (Lukasczyk and Fraser, 2022), a search-based test-generation tool, and improves upon it by leveraging LLM-generated test cases to expand the search space on reaching coverage plateaus. The available version of CodaMosa works with the deprecated code-davinci-002 model; we, therefore, updated the tool to work with GPT-3.5-turbo-instruct, the recommended replacement model for code-davinci-002 (OpenAI, 2024b).
@@
Figure 4. Line, branch, and method coverage achieved on Java SE and Java EE applications by aster (configured with different LLMs) and EvoSuite (GPT-4 run excluded for App X for confidentiality reasons).
4.3. Models
We selected six models for evaluation, including LLMs of different sizes (ranging from 8B to over 1T parameters), open-source and closed-source LLMs, and LLMs from different model families (GPT (OpenAI, 2024c), Llama-3 (Meta, 2024), and Granite (IBM, 2024)), considering general-purpose and code models. Table 1 provide details of the selected models. Our emphasis was on selecting open-source models along with one frontier model.
4.4. Datasets
Table 2 lists the datasets used in the evaluation. The Java dataset is split into Java Standard Edition (SE) and Enterprise Edition (EE) applications. For Java SE, we selected four Apache Commons projects: Commons CLI (commonscli, 2024), Commons Codec (commonscodec, 2024), Commons Compress (commonscompress, 2024), and Commons JXPath (commonsjxpath, 2024). For Java EE, we used three open-source applications (CargoTracker (cargotracker, 2024), DayTrader (daytrader, 2024), PetClinic (petclinic, 2024)), covering different Java frameworks and a proprietary enterprise application (called “App X” for confidentiality).
For Python, we started with 486 modules from the CodaMosa artifact (codamosaartifact, 2024). We ran CodaMosa on this dataset, but encountered crashes on 203 of the modules. We excluded those modules, resulting 286 modules from 20 projects.
4.5. Evaluation metrics
4.5.1. Code coverage and mutation scores
To measure code coverage, we used JaCoCo (EclEmma, [n. d.]) for Java and Coverage.py (Batchelder, [n. d.]) for Python. Because Coverage.py does not report method coverage, we implemented custom code for inferring method coverage from line coverage. For both Java and Python, we report data on line, branch, and method coverage. To compute mutation scores, we used PIT (PITEST, [n. d.]) for Java and mutmut (Hovmöller, [n. d.]) for Python. Both of these tools take an application and a test suite as inputs, generate mutants for the application by introducing small changes using mutation operators, and compute mutation scores for the test suite. We used the default set of mutation operators available in the tools.
4.5.2. Naturalness
For test cases, we consider naturalness to encompass different characteristics, such as (1) readability in terms of meaningfulness of test and variable names, (2) quality of test assertions, (3) meaningfulness of test sequences, (4) quality of input values, and (5) occurrences of test smells or anti-patterns. Our evaluation focuses on assessing characteristics 1 and 2 quantitatively; additionally, we conducted a developer survey, which provides developer perspectives on characteristics 1–4. For studying occurrences of test smells, we attempted to use a test-smell detection tool for Java, TSDetect (tsdetect, 2024; Peruma et al., 2020), but ran into numerous issues with the tool, such as identification of spurious test smells (false positives) and incorrect counts of test smells. We therefore chose to not use the tool, and perform our quantitative evaluation with a custom naturalness checker for characteristics (1) and (2) implemented using Tree-sitter (treesitter, 2024) and WALA (wala, 2024). We note that our metrics for assessing test assertion quality include some of the test smells detected by TSDetect.
Measuring assertion quality
  • Assertion ratio. This measures the percentage of lines of code with assertions in a test case.
  • Tests with no assertions. The percentage of test cases in a test file that have no assertions.
  • Tests with null assertions. For a test file, the percentage of tests with assertions that contain a null assertion.
Measuring meaningfulness of test method name
A unit test name should clearly capture the functionality being tested, by including the focal method name and the functionality being tested if necessary. For example, a test that checks whether command-line options are correctly populated when long option values are provided can be named test_getOptions_longOptions. However, tools such as EvoSuite and CodaMosa often generate non-descriptive test names (e.g., test01) that do not reflect the functionality being tested. We developed the following approach for measuring the meaningfulness of test names. First, for a given test file, we identify the focal class, similar to the Method2Test approach (Tufano et al., 2020). For Python, we skip this step because focal methods can exist inside or outside of classes, and resolving imports is more complex. Instead, we directly identify focal methods, assuming any function call could be a focal method. For Java, we examine all call sites in the test case, matching them with testable methods in the focal classes. This step results in a list of potential focal methods. We then check if one of these methods is mentioned in the test name. If it is, we assign a 50% score for the match. Next, we tokenize the remaining part of the test name after removing “test” keywords. We use camel-case and underscore splitting and generate all possible word combinations by merging them sequentially. For example, test_addOption_longArgs_throwsException is first matched with the focal method addOption, then the name is broken into long, args, and longargs. For exceptions, we match them separately with exceptions thrown in the test body. We calculate the closeness score using the Levenshtein distance by matching these tokens with all code identifiers.
Measuring meaningfulness of variable names in tests
Consider the below illustrative example, showing two test cases that have the same set of steps. Our premise is that the name of a variable of a data structure type (e.g., String, List) can be meaningful based on its context, whereas the meaningfulness of a variable name for a non-data structure type depends on both type and context.
Figure 5. Example of two tests with meaningfulness of method name and variable name.
For instance, a variable of type String named str is not meaningful as it is not capturing any information about the stored data. In contrast, variable BasicParser parser is meaningful as type name itself conveys meaning of the stored data. Additionally, a variable name’s meaningfulness can depend on the context in which it is used. For example, in the case of String[] arguments, the variable name depends on the formal parameter names of the method flatten(Options options, String[] arguments, boolean stopAtNonOption) whose returned value is stored in the variable. Based on this premise, we categorize variables into two groups. Then, depending on the group, we determine whether to match with the data type name, assignment context, and formal parameter names, or simply the assignment context and formal parameter names. Finally, we use the Levenshtein distance to compute the closeness score.
4.6. Experiment environment
@@
The experiments were conducted on cloud VMs, each equipped with a 48-core Intel(R) Xeon(R) Platinum 8260 processor; the RAM ranged from 128 to 384 GB. We used the OpenAI API (OpenAI, 2024a) to access the GPT models and an internal (proprietary) cloud service to access the other models. We used v1.2.0 of EvoSuite and our updated version of CodaMosa. We performed three runs of test generation with each model for both Java dataset and the Python dataset. We use temperature = 0.2, which has been used for code generation tasks (Ugare et al., 2024), and for generating more predictable outputs, with max token length set to 1024 for reducing cost. We use the same settings for all the models.
5. Related Work
@@
Compared to the previous works (Fraser and Arcuri, 2011; Arcuri, 2019; Pacheco et al., 2007; Pacheco and Ernst, 2007; Lukasczyk and Fraser, 2022; Tillmann and De Halleux, 2008; Sen et al., 2005; Godefroid et al., 2005; Tzoref-Brill et al., 2022) on conventional testing approaches, we leverage LLM to generate more natural unit tests while supporting multiple PLs. There have been several works that attempted to use LLMs for test generations (Wang et al., 2024b; Tufano et al., 2020; Vikram et al., 2023; Bareiß et al., 2022; Ryan et al., 2024; Schäfer et al., 2023). The closest works are done by ChatTester (Yuan et al., 2024) and ChatUnitTest (Xie et al., 2023) (arxiv), which have used the GPT model to generate test cases given the information related to the class under test and has shown that it performs competitively. However, this approach is specifically tied to GPT, and in many enterprise use cases, this approach may not work. Recent Pizzorno and Berger (Pizzorno and Berger, 2024) (arxiv), use LLM as a coverage augmenting approach starting with tests generated by CodaMosa. Compared to that work, aster (1) supports more PLs, (2) generates tests from scratch, (3) creates more natural tests (based on our automated analysis and developers’ survey) compared to CodaMosa, which the prior work is dependent on, and (4) generates tests with higher coverage (prior work: line: +7.3%, branch: +10.3%, aster: line: +37.1%, branch: 24.9% (with GPT-4, on which the prior work has been evaluated)) compared to CodaMosa.
6. Discussion
Our study identified several pros and cons of using LLMs for test generation. Although aster addresses some of them, there remains significant room for further research.
Supporting more PLs. Although LLMs are trained on various PLs, using them off-the-shelf may not always produce the best results. However, when combined with lighter-weight program analysis, their performance is often very competitive with, and sometimes surpasses, existing approaches. This reduced need for complex program analysis makes it easier to support more PLs.
Multiple frameworks and versions Another benefit of using LLMs for test generation is their ability to support multiple frameworks and PL versions. For instance, state-of-the-art tools such as EvoSuite do not support beyond Java 11–posing a significant challenge for practical usage. LLMs can play a significant role in bridging this gap.
Increasing naturalness of existing approaches. In this work, we developed an automated method to assess the naturalness of test cases and discovered that LLM-generated tests are notably more natural than those produced by existing tools or written by developers. We can leverage a similar approach to (1) enhance the naturalness of test cases generated by existing approaches and (2) guide LLMs to produce more natural code by incorporating this aspect as a pretraining objective.
Beyond unit testing. We evaluated LLMs’ capability in unit test generation. However, we believe that there is a significant potential of using LLMs to go beyond unit testing to other testing levels—module, integration, and system—for which there is limited support among the current test-generation tools, especially for white-box testing.
Affordable models. One of the main challenges of using LLMs is the associated cost. LLM-based approaches require substantial computing resources, making them less accessible for developers compared to existing methods. However, our study revealed some noteworthy findings: (1) unlike many other use cases, GPT-4 is not the most dominant model and (2) smaller models (e.g., Granite-34b and Llama3-8b) perform competitively. We believe that more effort should be directed towards developing quantized, test-generation-specific models that can run efficiently.
7. Threats to Validity
In this section, we identify the most relevant threats to validity and discuss how we mitigated them.
External threat. To address threat related to generalizability of aster, we extended support for (a) multiple PLs and (b) multiple models with varying sizes, modalities, and from different model families.

### v3

@@ III Our Approach
    Preprocessing for Python.  Unlike Java, Python’s more flexible structure, where a single file or module can contain a mix of functions, classes, and standalone code, lends itself better to module-scoped test generation. Therefore, for Python test generation, aster targets a module as a whole for test generation, including all classes, methods, and functions declared in the module. This module-level approach together with distinct language features of Python allows us to omit some preprocessing steps that are necessary for Java but irrelevant for Python. Because our approach provides the entire module to the LLM, instructing the LLMs about calls to private methods is unnecessary. Member visibility is Python is specified via naming conventions (names beginning with single underscore for protected members and double underscores for private members) and there is no strict encapsulation as these members can be accessed via name mangling. Python’s properties feature (using @property decorators) also lets private members to be accessed directly. Thus, identification of accessors is also unnecessary for Python test generation. In terms of relevant constructors, the constructor definitions in the focal module are already available to the LLM. Additionally, we add constructors for all imported modules to the LLM prompt. Also, we found that, for Python, adding a few examples of tests helps LLMs produce more predictable output. For Java, this was not necessary, but investigation of RAG-based approaches for incorporating in-context learning for test generation is an interesting future research direction.
Postprocessing
Despite using rich context gathered through static analysis, LLMs can still generate code that: (1) does not adhere to PL syntax (non-parsable), (2) does not adhere to the testing framework, (3) does not have sufficient details such as imports, package name, class name, folder location (for Java tests, it is essential to follow the source directory structure), (4) has compilation issues, or (5) runs but fails due to assertion or runtime failures. During postprocessing, we perform several fixing steps to remediate these issues. Additionally, we perform coverage augmentation to increase the coverage of the initial set of tests generated by aster.
III-1 Output sanitization
First, we sanitize the generated code for extraneous content (e.g., natural language text) and ensure the syntactic correctness of the generated test cases. We also gather the required imports by static analysis. For instance, we look into the focal class and its superclasses to gather the used imports. We add all application classes as imports because, often, test cases refer to other application classes that may not be imported in the focal class or its superclasses. We also add a few imports related to the testing framework. Any unused imports added in this process is removed at the end of test generation. Next, for Java, we augment the build file to add the testing framework dependencies, if required, and identify the Java and JUnit versions from the build file, which are used for generating test cases. Third, we check the names of the test and setup/teardown methods to avoid duplicate content as the LLM may generate test cases without paying attention to details. Fourth, we add the test framework details, such as annotations, timeout values, etc., and also add exception information. Finally, we form the test class for each focal class and store it to desired location with all essential details. These steps are very crucial as, even with the best LLM, without code sanitization, a significant proportion of the generated tests may end up with compilation or runtime errors.
III-2 Error remediation
After sanitizing the code, we perform compilation and runtime checks. For each type of error, we first apply a set of rule-based fixes. For instance, we try to resolve errors due to naming clash of methods, variables, etc.; we attempt to fix assertion failures where the expected and actual values do not match. Also, we localize the error, decompose tests with multiple assertions into multiple tests with single assertions, and fix them. At the end of test generation, we merge the tests back to reduce the number of test cases. Then, for the rest of the errors, we gather all the context required for fixing the error, such as error message from compiler feedback, erroneous line, and context surrounding the focal method and re-prompt LLM for fixing the test cases (depicted in fig. 4- 3). We sanitize the fixed code and add them to the test suite.
III-3 Increasing coverage of generated tests
@@
Error-free test cases are executed to measure code coverage, which is further analyzed to identify uncovered code lines and update the LLM prompt template (fig. 4) to guide the LLM to generate test cases targeting the uncovered code lines. We also add the context gathered in the preprocessing phase to the prompt.
@@
Postprocessing for Python. For Python, we use Pylint [30], for identifying compilation and parsing errors. Specifically, aster focuses on the error category of Pylint warnings and uses this information to guide LLMs in fixing compilation and parsing issues. Then, tests are executed using Pytest [31] and the output is used to provide feedback to the LLMs in case of failures. Finally, Coverage.py [32] is used to measure coverage and perform the coverage augmentation.
TABLE I: Models used in the evaluation.
@@
IV Experiment Setup
A. Research Questions: Our evaluation focuses on three research questions.
RQ1:
@@ IV Experiment Setup
TABLE II: Java and Python datasets used in the evaluation.
@@
B. Baseline test-generation tools: We evaluated aster against two state-of-the-art unit test generators. For Java, we used EvoSuite [22], specifically Release 1.2.0. For Python, we selected CodaMosa and used the latest version from its repository [23]. CodaMosa is built on Pynguin [13], a search-based test-generation tool, and improves upon it by leveraging LLM-generated test cases to expand the search space on reaching coverage plateaus. The available version of CodaMosa works with the deprecated code-davinci-002 model; we, therefore, updated the tool to work with GPT-3.5-turbo-instruct, the recommended replacement model for code-davinci-002 [33].
C. Models: We selected six models for evaluation, including LLMs of different sizes (ranging from 8B to over 1T parameters), open-source and closed-source LLMs, and LLMs from different model families (GPT [34], Llama-3 [35], and Granite [36]), considering general-purpose and code models. Table I provide details of the selected models.
D. Datasets: Table II lists the datasets used in the evaluation. The Java dataset is split into Java Standard Edition (SE) and Enterprise Edition (EE) applications. For Java SE, we selected four Apache Commons projects: Commons CLI [28], Commons Codec [37], Commons Compress [38], and Commons JXPath [29]. For Java EE, we used three open-source applications (CargoTracker [39], DayTrader [40], PetClinic [41]), covering different Java frameworks and a proprietary enterprise application (called “App X” for confidentiality).
For Python, we started with 486 modules from the CodaMosa artifact [42]. We ran CodaMosa on this dataset, but encountered crashes on 203 of the modules. We excluded those modules, resulting 283 modules from 20 projects. We went through these failed modules individually to check whether they were indeed failures.
E. Evaluation metrics:
Code coverage: We used JaCoCo [43] for Java and Coverage.py [32] for Python to measure coverage. Because Coverage.py does not report method coverage, we implemented custom code for inferring method coverage from line coverage.
Naturalness: For test cases, we consider naturalness to encompass different characteristics, such as (1) readability in terms of meaningfulness of test and variable names, (2) quality of test assertions, (3) meaningfulness of test sequences, (4) quality of input values, and (5) occurrences of test smells or anti-patterns. Our evaluation focuses on assessing characteristics 1 and 2 quantitatively; additionally, we conducted a large-scale developer survey, which provides developer perspectives on characteristics 1–4. For studying occurrences of test smells, we attempted to use a test-smell detection tool for Java, TSDetect [44, 45], but ran into numerous issues with the tool, such as identification of spurious test smells (false positives) and incorrect counts of test smells. We therefore chose to not use the tool, and perform our quantitative evaluation with a custom naturalness checker for characteristics (1) and (2) implemented using Tree-sitter [46] and WALA [47]. We note that our metrics for assessing test assertion quality include some of the test smells detected by TSDetect.
Measuring assertion quality
  • Assertion ratio. This measures the percentage of lines of code with assertions in a test case. Generally, tests with too many assertions can be brittle and break easily.
  • Tests with no assertions. The percentage of test cases in a test file that have no assertions. Tests without assertions are considered an anti-pattern as they do not check for expected behavior and can only detect crashing failures.
  • Tests with null assertions. For a test file, the percentage of tests with assertions that contain a null assertion, which is considered an anti-pattern.
Measuring meaningfulness of test method name
A unit test name should clearly capture the functionality being tested, by including the focal method name and the functionality being tested if necessary. For example, a test that checks whether command-line options are correctly populated when long option values are provided can be named test_getOptions_longOptions. However, tools such as EvoSuite and CodaMosa often generate non-descriptive test names (e.g., test01) that do not reflect the functionality being tested. We developed the following approach for measuring the meaningfulness of test names. First, for a given test file, we identify the focal class, similar to the Method2Test approach [48]. For Python, we skip this step because focal methods can exist inside or outside of classes, and resolving imports is more complex. Instead, we directly identify focal methods, assuming any function call could be a focal method. For Java, we examine all call sites in the test case, matching them with testable methods in the focal classes. This step results in a list of potential focal methods. We then check if one of these methods is mentioned in the test name. If it is, we assign a 50% score for the match. Next, we tokenize the remaining part of the test name after removing “test” keywords. We use camel-case and underscore splitting and generate all possible word combinations by merging them sequentially. For example, test_addOption_longArgs_throwsException is first matched with the focal method addOption, then the name is broken into long, args, and longArgs. For exceptions, we match them separately with thrown exceptions. We calculate the closeness score using the Levenshtein distance by matching tokens with code identifiers.
Measuring meaningfulness of variable names in tests
Consider the following example, showing two test cases that have the same set of steps. Our premise is that the name of a variable of a data structure type (e.g., String, List) can be meaningful based on its context, whereas the meaningfulness of a variable name for a non-data structure type depends on both type and context.
For instance, a variable of type String named str is not meaningful as it is not capturing any information about the stored data. In contrast, variable BasicParser parser is meaningful as the type name itself conveys meaning of the stored data. Additionally, a variable name’s meaningfulness can depend on the context in which it is used. For example, in the case of String[] arguments, the variable name depends on the formal parameter names of the method flatten(Options options, String[] arguments, boolean stopAtNonOption) whose returned value is stored in the variable. Based on this premise, we categorize variables into two groups. Then, depending on the group, we determine whether to match with the data type name, assignment context, and formal parameter names, or simply the assignment context and formal parameter names. Finally, we use the Levenshtein distance to compute the closeness.
@@ E. Evaluation metrics:
Figure 5: Line, branch, and method coverage achieved on Java SE and Java EE applications by aster (configured with different LLMs) and EvoSuite (GPT-4 run excluded for App X for confidentiality reasons).
@@
F. Experiment environment: The experiments were conducted on cloud VMs, each equipped with a 48-core Intel(R) Xeon(R) Platinum 8260 processor; the RAM ranged from 128 to 384 GB. We used the OpenAI API [49] to access the GPT models and an internal (proprietary) cloud service to access the other models. We used v1.2.0 of EvoSuite and our updated version of CodaMosa. We performed three runs of test generation with each model for both Java dataset and the Python dataset. We use temperature = 0.2, which has been used for code generation tasks [50], and for generating more predictable outputs, with max token length set to 1024 for reducing cost.
V Lessons Learned
Supporting more PLs and frameworks. LLMs, when combined with lightweight program analysis, offer versatile support for multiple PLs and frameworks. The aster pipeline demonstrates this adaptability, often matching or surpassing existing methods in test generation across various languages. To extend support, developers can utilize tools like Tree-sitter [46], CodeQL [54], and WALA [47], addressing the crucial need for multi-language and multi-framework support in both legacy and continuous development enterprise contexts.
Generating more natural test cases. Developers prefer naturalness in test cases, including intuitive test names, variable names, and high-quality assertions. LLMs, trained on developer-written code, inherently produce more natural output compared to existing ATGs. While LLMs can make mistakes, combining them with program analysis yields high-quality test cases in terms of both coverage and naturalness. Potential research directions in this area include enhancing the naturalness of conventionally generated tests and incorporating naturalness as a pretraining objective for LLMs.
Affordability and privacy. LLM-based test generation faces affordability and privacy challenges in enterprise settings. Our study shows smaller models such as Granite-34b and Llama3-8b perform competitively, suggesting cost-effective alternatives to GPT-4. Privacy concerns necessitate on-premise solutions, with developers preferring models hosted internally or on local workstations. Future research should focus on developing efficient, quantized models specific to test generation, and reducing LLM calls through deterministic approaches.
VI Related Work
@@
Compared to the previous work [11, 55, 14, 56, 13, 7, 5, 4, 57] on conventional test-generation approaches, we leverage LLMs to generate more natural unit tests while supporting multiple PLs. Recently, there have been several attempts at employing LLMs for test generation [58, 48, 59, 60, 61, 62, 63]. AthenaTest [48] developed a fine-tuning pipeline for test generation. Vikram et al. [59] leveraged LLMs to support property-based testing. Ryan et al. [61] adopted symbolic analysis to boost coverage and Bareiß et al. [60] incorporated few-shot examples to guide test generation. TestPilot [62] synthesizes unit test cases for JavaScript and TypeScript by gathering usage documentation and API functions. ChatTester [64] and ChatUnitTest [65] used the GPT model to generate test cases given the information related to the class under test and have shown that it performs competitively. However, this approach is specifically tied to GPT, and in many enterprise use cases, this approach may not work. Recent Pizzorno and Berger [66] (arxiv), use LLM as a coverage-augmenting approach starting with tests generated by CodaMosa. Compared to that work, aster supports more PLs, generates tests from scratch, creates more natural tests than CodaMosa, which the prior work is dependent on, and generates tests with higher coverage than CodaMosa (prior work: line: +7.3%, branch: +10.3%, aster: line: +37.1%, branch: +24.9% with GPT-4, on which the prior work was evaluated).
Compared to the studies done in different industry settings [67, 25, 68, 62], our work provides several new findings and also confirms various important lessons learned in the prior work. For instance, the acceptance rate of the LLM-generated tests reported by the work done at Meta [25] and our work are very similar (73% at Meta and $>$70% in our case), which shows that developers overall prefer such test cases. Moreover, both studies identify test naturalness as a key factor in developers’ acceptance criteria. Another common finding of these studies is that using static and dynamic analysis help increase the performance of the LLMs. Our work provides several new findings, related to generation of tests with API mocking, multi-language test generation, performance of models of different sizes/families, and a rigorous study of test naturalness including a developer survey.
VII Threats to Validity